Three times the smallest of three consecutive integers is six more than twice the largest. Find the integers.
Found 2 solutions by fractalier, stanbon:
Answer by fractalier(6550)

Call the integers x, x+1, and x+2. Then we have
3x = 2(x+2) + 6
3x = 2x + 4 + 6 = 2x + 10
x = 10
The integers are 10, 11, and 12.

Answer by stanbon(75887)

Three times the smallest of three consecutive integers is six more than twice the largest. Find the integers.
-----
1st: x-1
2nd: x
3rd: x+1
------
Equation:
3(x-1) = 2(x+1)+6
-----
3x - 3 = 2x + 8
x = 11
----
x-1 = 10
x+1 = 12
--------------
Cheers,
Stan H.
-------------